to calculate velocity of a car accelerating from rest in straight line. Equation is v(t) = A (1-e ^ -t/t max speed) the ^ means to the power of. v(t) is the instantaneous velocity of the car (m/s), t is the time in seconds, tmaxspeed is the time to reach the maximum speed in seconds and A is a constant. 1. Create a graph of acceleration vs time? 2. Apply the mathematical model which is a car with a t(0-28m/s) of 2.6s, a t(400m) of 10.46s in order to calculate the value of coefficient, maximum velocity and maximum acceleration?
1
Expert's answer
2020-01-10T08:34:21-0500
"\text v(t)=A(1-e^{-{t \over t_{max\_ speed}}})"

"\text a(t)=\text v'(t)={A \over t_{max\_ speed}}e^{-{t \over t_{max\_speed}}}"
"\text v(t)=A(1-e^{-{t \over t_{max\_ speed}}})"

"\text s(t)=\int \text v(t) dt=\int A(1-e^{-{t \over t_{max\_ speed}}})dt="

"=A(t+t_{max\_speed}\cdot e^{-{t \over t_{max\_ speed}}})+C"

"\text s(0)=0=>C=-A\cdot t_{max\_speed}"

"\text s(t)=A(t-t_{max\_speed}+t_{max\_speed}\cdot e^{-{t \over t_{max\_ speed}}})"

Given that

"\text v(2.6)=A(1-e^{-{2.6 \over t_{max\_ speed}}})=28"

"\text s(10.46)=""A(10.46-t_{max\_speed}+t_{max\_speed}\cdot e^{-{10.46 \over t_{max\_ speed}}})=400"


"{10.46-t_{max\_speed}+t_{max\_speed}\cdot e^{-{10.46 \over t_{max\_ speed}}} \over 1-e^{-{2.6 \over t_{max\_ speed}}}}={400 \over 28}"

"t_{max\_speed}\approx5.834\ s" or "t_{max\_speed}\approx28.424 \ s"

"A\approx77.864\ m/s" or "A\approx320.318\ m/s"


"\text v_{max}=A(1-e^{-{ t_{max\_ speed} \over t_{max\_ speed}}})=A(1-e^{-1})"


"\text a_{max}={A \over t_{max\_ speed}}e^{-{t_{max\_speed} \over t_{max\_speed}}}={Ae^{-1} \over t_{max\_ speed}}"

"t_{max\_speed}\approx5.834\ s,A\approx77.864\ m/s"


"\text v_{max}=77.864(1-e^{-1})\ m/s\approx49.219\ m/s"

"\text a_{max}={77.864\ m/s\cdot e^{-1} \over 5.834\ s}\approx4.910\ m/s^2"

Or

"t_{max\_speed}\approx28.424\ s,A\approx320.318\ m/s"


"\text v_{max}=320.318(1-e^{-1})\ m/s\approx202.480\ m/s"

"\text a_{max}={320.318\ m/s\cdot e^{-1} \over 28.424\ s}\approx4.146\ m/s^2"

"202.480\ m/s\approx729\ km/h"

I think, that the car cannot have such speed.